OpenStudy (anonymous):

Use Cramer's rule to determine if the system is inconsistent system or contains dependent equations. 2x + 7 = 8 6x + 3y = 24 Answer Options: A. system is inconsistent B. system contains dependent equations Thanks!

OpenStudy (anonymous):

B

OpenStudy (anonymous):

So your equations are 2x + 0y = 1 6x + 3y = 24 x=1/2 3 + 3 y =24 3y = 21 y =7

OpenStudy (anonymous):

Oh thanks! That was my thought, but I wasn't sure.

OpenStudy (anonymous):

No problem(: